OpenStudy (anonymous):

Complete the table and choose the solution for each given value of x y= -x2 +2 when x =-2 not understanding how to do it

OpenStudy (anonymous):

is "-x2" supposed to mean x squared?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

if it does then -2 times -2 equals positive 4 plus the 2 and the answer is 6

OpenStudy (anonymous):

it took me a few times to understand the question but if you just read slowly you can understand it

OpenStudy (anonymous):

But the answer choices are y = -(2)^ + 2=-2;(-2,-2) that's one of them

OpenStudy (anonymous):

you have a y but no x, and what is the "-(2)" to the power of?

OpenStudy (anonymous):

X =-2

OpenStudy (anonymous):

if this is the same equation as your original then, y=-2^2+2 put in the y value and it's, -2= -2^2 + 2

OpenStudy (anonymous):

solve and it does not work because the x has to equal out to the y value when you finish simplifying

OpenStudy (anonymous):

if you simplified this then it goes to -2=6 not -2=-2 see what I mean?

OpenStudy (anonymous):

do you have any other answer choices?

OpenStudy (anonymous):

Yes I see

OpenStudy (anonymous):

do you have any other answer choices?

OpenStudy (anonymous):

Y=(-2^) +2=2;(-2,2) Y = -(2)^+2=-2;(2,-2)

OpenStudy (anonymous):

y= -(-2)^-2=2;(-2,-2)

OpenStudy (anonymous):

you forgot to put the power 2

OpenStudy (anonymous):

Oops

OpenStudy (anonymous):

weel we are getting close but still it comes out to -2=2

OpenStudy (anonymous):

oops *well

OpenStudy (anonymous):

y= -(-2)^-2=2;(-2,-2) is that other negative sigh out side of th () supposed to be there?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Y = -(2)^+2=-2;(2,-2) now that I look back this could be the solution if the negative sign is supposed to be on the outside of the ()

OpenStudy (anonymous):

if it is then it's -2=-()^2 +2: 2x2=4 and since the negative is on the outside the 4 become negative: add the -4 and positive 2 to get -2=-2

OpenStudy (anonymous):

oops -2=-()^2 +2 this is supposed to look like -2=-(2)^2 +2 sorry

OpenStudy (anonymous):

but yes -2=-(2)^2 +2 should be your anser I believe it's this one: Y = -(2)^2+2=-2;(2,-2)

OpenStudy (anonymous):

How about when x=-1

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

is there an equation that comes with that or is it the same one we just did?

OpenStudy (anonymous):

I have four more like the first one four more problems math is my worse subject

OpenStudy (anonymous):

well I need the equation and y value to do them

OpenStudy (anonymous):

what is the next question?

OpenStudy (anonymous):

It says choose the solution for each given value of x. Y= -x2 +2 Then it says when x = -1 When x= 0 When x = 1 When x = 2

OpenStudy (anonymous):

oh ok so we just switch the x value

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so I guess we are solving for y?

OpenStudy (anonymous):

I guess I don't know how to do it. Thanks for the help I really stink at math

OpenStudy (anonymous):

Y= -x2 +2 plug in the known values and it looks like y= -1^2 + 2. Now 1 times 1 equals 1 so 1 + 2 equals 3 this means y=3

OpenStudy (anonymous):

you keep forgetting the arrow for -x^2

OpenStudy (anonymous):

Oops sorry

OpenStudy (anonymous):

or did you copy and paste the question like that?

OpenStudy (anonymous):

No I forgot it

OpenStudy (anonymous):

oh okay, anything else, diod you do x=0?

OpenStudy (anonymous):

*did sorry

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok any other questions?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

ok see ya

OpenStudy (anonymous):

Thanks